OpenStudy (anonymous):

The cylinder below has a volume of 1,538.6 cubic inches. What is the radius of the cylinder? (Use 3.14 for π.)

OpenStudy (anonymous):



OpenStudy (lgbasallote):

\[Volume = \pi r^2 h\] \[1538.6 = \pi r^2(10)\] \[153.86 = \pi r^2\] \[\frac{153.86}{\pi} = r^2\] \[\sqrt{\frac{153.86}{\pi}} = r\] got it?